(a) The graph of y=f(x) is shown. Draw the graph of y=f(-x).
(b) The graph of y=g(x) is shown. Draw the graph of y=-g(x).

Answer by ikleyn(52784)

.
(a) The graph of y=f(x) is shown. Draw the graph of y=-f(x).
(b) The graph of y=g(x) is shown. Draw the graph of y=g(-x).
~~~~~~~~~~~~~~~~~~~~

(a)  The graph y = -f(x) is the mirror reflection of the graph y = f(x) 
     relative x-axis.



(b)  The graph y = f(-x) is the mirror reflection of the graph y = f(x) 
     relative y-axis.

Solved.